# Analyzing the Effect of Linguistic Similarity on Cross-Lingual Transfer: Tasks and Experimental Setups Matter

### v2

@@ Abstract
Cross-lingual transfer is a popular approach to increase the amount of training data for NLP tasks in a low-resource context. However, the best strategy to decide which cross-lingual data to include is unclear. Prior research often focuses on a small set of languages from a few language families and/or a single task. It is still an open question how these findings extend to a wider variety of languages and tasks. In this work, we analyze cross-lingual transfer for 266 languages from a wide variety of language families. Moreover, we include three popular NLP tasks: POS tagging, dependency parsing, and topic classification. Our findings indicate that the effect of linguistic similarity on transfer performance depends on a range of factors: the NLP task, the (mono- or multilingual) input representations, and the definition of linguistic similarity.
@@ 1 Introduction
For many of the world’s languages, the available data to train machine learning models is scarce. If data is available, it is often only enough for a validation set, raising the question of how to select the training set. Two approaches are intuitive: (i) based on linguistic similarity measures: find the training data in a language that is linguistically as close as possible to the target language, where “linguistically close” is defined by one or multiple linguistic similarity measures, or (ii) based on dataset dependent measures: find a dataset in another language that is similar to the target dataset, e.g., in terms of high word or $n$-gram overlap. Naturally, these two types of approaches are not mutually exclusive.
@@
In this work, we contribute to these questions by analyzing transfer between $266$ different languages from $33$ language families (Figure 1) in three different tasks: (i) part-of-speech (POS) tagging, (ii) dependency parsing, and (iii) topic classification. In choosing our tasks, we are motivated by the following considerations: (i) availability of datasets: we select tasks for which we can find data for a multitude of languages from different language families, and (ii) types of tasks: we include two word-level grammatical tasks (i.e., POS tagging and dependency parsing) and one sentence-level topic identification task. To allow for a clean analysis, we opt for a zero-shot approach, in line with prior work de Vries et al. (2022) – we train our models on the task in one particular source language, and evaluate on the target language without additional training or fine-tuning in the target language.
@@ 2 Related Work
Table 1: Related work focusing on zero-shot transfer between many languages or on many tasks. Tasks: D=dependency parsing, E=entity linking, I=natural language inference, M=machine translation, N=named entity recognition, P=part-of-speech tagging, Q=question answering, S=sentence retrieval, T=topic classification.
@@ 3 Methodology
We run transfer experiments on two word-level grammatical tasks with word-level annotations (§3.1) and a sentence-level topic classification task (§3.2). Of the 266 languages in our experiments, 55 training and 84 test languages are shared between all tasks. Appendix §A contains details on all languages and datasets in this study. §3.3 introduces our similarity measures.
@@ 3.1 Grammatical tasks
We use the UDPipe 2 models (Straka, 2018, 2023) that were trained on UD 2.12 data. Each model is trained on a single treebank. The models have so far only been systematically evaluated on their within-treebank performance, but not cross-lingually. UDPipe 2 combines monolingual character and word embeddings [^2] The embeddings are trained on the available training treebanks for a given language. with multilingual embeddings derived from mBERT Devlin et al. (2019). The models are trained jointly for POS tagging, dependency parsing, lemmatization and morphological feature prediction. We only evaluate on the first two, as many treebanks do not have gold standard labels for the others. UDPipe 2 post-processes the predicted dependencies to ensure that each sentence includes a root dependency that all other nodes (in)directly depend on.
@@ 3.2 Topic classification
We use multi-layer perceptrons (MLPs) for topic identification, similar to the baseline models by Adelani et al. (2024). This lightweight architecture allows training and evaluating many models without prohibitive time or energy investments, and results in evaluation scores that are close to the performance of base-sized transformers like XLM-R Conneau et al. (2020)E, Table 9). We use the scikit-learn implementation (Pedregosa et al., 2011) and conduct hyperparameter tuning on a subset of the languages (details in Appendix §D). We compare different ways of representing the input data:
@@ 3.3 Similarity measures
We determine whether two languages are related (and how closely) with the help of Glottolog’s Hammarström et al. (2024) phylogenetic trees. For each language, we retrieve the path from its family root node to the language node (details in §C.1). The relatedness of two languages (gen) is the ratio of shared nodes along these paths. [^4] Because different trees/branches have different depths (owing to different family sizes and/or different levels of documentation), comparisons of scores can only serve as a proxy for the degree of phylogenetic relatedness. This is also noted in the metadata for lang2vec, which calculates genetic distances in a similar way: http://www.cs.cmu.edu/~dmortens/projects/7_project/.
@@
We use lang2vec’s location information (each language is represented as a vector of distances to a number of points on the Earth’s surface) and calculate the Euclidean distances between language vectors. We define geographic proximity (geo) as 1 minus the distance. [^5] We do not use lang2vec’s precomputed distances, as the package erroneously returns the maximum distance for some languages, regardless of the comparison.
@@
Correlations between similarity measures can influence how importance is assigned to different measures in regression analyses. Table 7 in §C.2 shows how the different similarity measures are correlated with each other. Lexical similarity and phylogenetic relatedness are highly correlated ($r=0.87$). Additionally, the two grammar-based similarity measures (gb and syn) are moderately strongly correlated ($r=0.61$), as are gb and lexical/phylogenetic similarity ($r=0.57$, $0.56$), and word overlap in UD and lexical similarity ($r=0.59$).
@@ 4 Supporting Results
Here we present the results on which the main analyses in §5 build. In §4.1 we present the POS tagging, parsing and topic classification scores within and across languages. In §4.2 we compare how (dis)similar transfer patterns are across tasks and experiments, which helps contextualize the importance of different similarity measures for different experiments and tasks in §5.1.
We focus on the language level for most of our analyses. For UD, which has nearly twice as many datasets as languages, we find that trends are very stable across datasets of the same language (§E.2).
@@
As expected, the within-language performances are much higher than the cross-lingual performances. Table 2 shows the mean scores within and across languages. Heatmaps with individual scores can be found in §E.1. Table 9 in §E additionally provides dataset-level scores and comparisons to scores reported in related work. Most notably, we find that our POS tagging results are similar but not identical to the ones that de Vries et al. (2022) obtained on UD v2.8 with fine-tuned XLM-R models for 65 training and 105 test languages. When we select the v2.8 languages and use language-level averages, our results are similar (Pearson’s $r$ and Spearman’s $\rho=0.73$, $p<0.0001$) – while the model choice (incl. input representations) matters for the resulting patterns, many transfer trends are similar.
Within-lang. Cross-ling. POS (accuracy) 93.4 6.8 43.0 20.2 Parsing (UAS) 84.8 10.1 36.8 18.8 Parsing (LAS) 78.4 13.7 20.6 17.0 topics-base (acc.) 66.7 14.1 20.1 8.8 topics-translit (acc.) 66.7 10.8 22.5 8.2 topics-mbert (acc.) 58.2 19.9 42.6 20.4
Table 2: Mean scores (in %) of evaluations within vs. across languages (standard deviations in subscripts).
@@
Although topics-base and topics-translit achieve the same within-language accuracy, topics-translit performs slightly better cross-lingually, having the advantage of higher $n$-gram overlap between datasets. [^7] We further analyze the effect of writing systems in §E.3. In our correlation analysis (§5.1), (lack of) character overlap serves as a proxy for writing system differences. In cross-lingual evaluations, topics-mbert benefits from its multilingual input embeddings: its accuracy is about twice as high (42.6%) as for the monolingual models. Generally, transfer works well if both training and test languages are in mBERT’s pre-training data (or closely related to a language that is): The average accuracy is 68.3% if both languages were included in mBERT’s pre-training data, 46.5% if only the target language was included, 36.1% if only the source language was included, and 33.2% if neither was included. This is consistent with results by Adelani et al. (2024), who fine-tuned four XLM-Rlarge models on one high-resource language each and found their transfer results to be very similar to each other.
@@
Figure 2: Different experiments produce different transfer patterns. NLP transfer results for all combinations of training (columns) and test languages (rows). The darker a cell, the higher the score. The three heatmaps for the grammatical tasks are sorted in the same order, and the three heatmaps with the topic classification results are sorted in the same order. The darker diagonal shows the within-language scores. Large, labelled heatmaps are in Appendix §E.
@@ 4.2 Comparing transfer patterns across tasks
In Figure 2, we plot heatmaps of the transfer results for the different experiments. Darker colors correspond to higher scores. To save space, we omit the language labels (which can be found in §E.1). We compute correlations between the results, which we summarize below (details in §F.1).
@@ 5.1 How do the similarity measures correlate with the tasks and input representations?
We calculate the correlations between the similarity measures and the NLP task results. To account for differences between test sets, we calculate correlations (Pearson’s r) between similarity measures and transfer results on a test language level. We compare the performance of different NLP systems on the same test language, but we do not compare the performance of a single system across multiple test sets. This decision is motivated especially by the challenges of comparing parsing performance across treebanks. Sentence length influences parsing difficulty as it determines the available search space (cf. McDonald and Nivre, 2011, Choi et al., 2015), and morphological differences between languages hamper the comparability of parsers Nivre and Fang (2017).
We analyze overall correlation scores by averaging across test languages. [^9] We treat correlation coefficients with p-values of at least 0.05 as 0, and exclude items where we could not calculate similarity scores due to missing linguistic information. We use language-level averages so that those languages with multiple training and test datasets do not have artificially high correlations, and languages with multiple test datasets do not have greater influence on the overall averages. However, if we instead consider treebank-level correlations, we observe that test sets that belong to the same language show very similar correlation patterns and scores (§E.2).
The correlations between task results and similarity measures vary across our experiments. Figure 3 shows the correlations, which we summarize below. Unaggregated correlations and 95% confidence intervals of the mean scores can be found in §F.3. These unaggregated correlations show that, while there are some outliers, the overall trends we report below accurately reflect the language-level trends.
@@
We observe some common trends across experiments: Training dataset size does not matter much: it is the same across all topic classification experiments, and the correlations between training set size and performance in the grammatical tasks are close to zero. The phonological and phonetic similarity measures (pho, inv) also generally have low correlation scores. [^10] Philippy et al. (2023) find that phonological similarity is more important for token- than sentence-level tasks. This is different in our study, where the correlations with inv are higher for the $n$-gram-based topic classification models. We assume that this is due to phonological similarity being moderately correlated with string similarity measures like trigram overlap (§C.2), which is in turn strongly correlated with the topic classification results.
The strongest predictor for parsing performance is syntactic similarity (syn) as determined by lang2vec ($r$avg_syn$=0.57$), followed by the similarity of Grambank features (gb; $r$avg_gb$=0.42$). These correlations are even stronger when we only consider the test languages that mBERT was pretrained on ($r$avg_syn$=0.69$ and $r$avg_gb$=0.60$, respectively; Appendix §F.3).
@@
The transfer results cannot be predicted by any one factor alone. We fit a linear mixed effects model for each experiment, with the NLP score as the dependent variable and the source and target languages as random effects. The fixed effects are the similarity scores, whether the training and test data use the same writing systems, and whether the test language was included in mBERT’s pre-training data. We use R R Core Team (2024) and the lme4 package Bates et al. (2015). This analysis involves a reduced set of languages, as for many language pairs in our data at least one similarity metric is not defined and we do not perform any data imputation. The analysis shows similar trends to the correlations described above. Additionally, for each experiment, most of the variables in our analyses are significant predictors of the transfer score (details in §F.2). This even sometimes applies to measures that capture similarity at the same linguistic level: e.g., both grammatical similarity measures (syn, gb) are independently among the highest predictors for parsing performance. The finding that multiple measures contribute to predicting the transfer score confirms previous work Lin et al. (2019); de Vries et al. (2022).

### v3

@@ Abstract
Cross-lingual transfer is a popular approach to increase the amount of training data for NLP tasks in a low-resource context. However, the best strategy to decide which cross-lingual data to include is unclear. Prior research often focuses on a small set of languages from a few language families and/or a single task. It is still an open question how these findings extend to a wider variety of languages and tasks. In this work, we analyze cross-lingual transfer for 263 languages from a wide variety of language families. Moreover, we include three popular NLP tasks: POS tagging, dependency parsing, and topic classification. Our findings indicate that the effect of linguistic similarity on transfer performance depends on a range of factors: the NLP task, the (mono- or multilingual) input representations, and the definition of linguistic similarity.
@@ 1 Introduction
For many of the world’s languages, the available data to train natural language processing (NLP) models is scarce. If data is available, it is often only enough for an evaluation set, raising the question of how to select the training set. Two approaches are intuitive: (i) based on linguistic similarity measures: find the training data in a language that is linguistically as close as possible to the target language, where “linguistically close” is defined by one or multiple linguistic similarity measures, or (ii) based on dataset dependent measures: find a dataset in another language that is similar to the target dataset, e.g., in terms of high word or $n$-gram overlap. Naturally, these two types of approaches are not mutually exclusive.
@@
In this work, we contribute to these questions by analyzing transfer between 263 different languages from 33 language families (Figure 1) in three different tasks: (i) part-of-speech (POS) tagging, (ii) dependency parsing, and (iii) topic classification. In choosing our tasks, we are motivated by the following considerations: (i) availability of datasets: we select tasks for which we can find data for a multitude of languages from different language families, and (ii) types of tasks: we include two word-level grammatical tasks (i.e., POS tagging and dependency parsing) and one sentence-level topic identification task. To allow for a clean analysis, we opt for a zero-shot approach, in line with prior work de Vries et al. (2022) – we train our models on the task in one particular source language, and evaluate on the target language without additional training or fine-tuning in the target language.
@@ 2 Related Work
Table 1: Related work focusing on zero-shot transfer between many languages or on many tasks. Tasks: D=dependency parsing, E=entity linking, F=natural language inference, I=intent classification, M=machine translation, N=named entity recognition, P=part-of-speech tagging, Q=question answering, S=sentence retrieval, T=topic classification.
@@ 3 Methodology
We run transfer experiments on two word-level grammatical tasks with word-level annotations (§3.1) and a sentence-level topic classification task (§3.2). Of the 263 languages in our experiments, 55 training and 84 test languages are shared between all tasks. Appendix §A contains details on all languages and datasets in this study. Section §3.3 introduces our similarity measures.
@@ 3.1 Grammatical tasks
We use the UDPipe 2 models (Straka, 2018, 2023) that were trained on UD 2.12 data. Each model is trained on a single treebank. The models have so far only been systematically evaluated on their within-treebank performance, but not cross-lingually. UDPipe 2 combines monolingual character and word embeddings with multilingual embeddings derived from mBERT Devlin et al. (2019). The models are trained jointly for POS tagging, dependency parsing, lemmatization and morphological feature prediction. We only evaluate on the first two, as many treebanks do not have gold standard labels for the others. UDPipe 2 post-processes the predicted dependencies to ensure that each sentence includes a root dependency that all other nodes (in)directly depend on.
@@ 3.2 Topic classification
We use multi-layer perceptrons (MLPs) for topic identification, similar to the baseline models by Adelani et al. (2024). This lightweight architecture allows training and evaluating many models without prohibitive time or energy investments, and results in evaluation scores that are close to the performance of base-sized transformers like XLM-R Conneau et al. (2020)D, Table 2). We use the scikit-learn implementation (Pedregosa et al., 2011) and conduct hyperparameter tuning on a subset of the languages (details in Appendix §C). We compare different ways of representing the input data:
@@ 3.3 Similarity measures
We determine whether two languages are related (and how closely) with the help of Glottolog’s Hammarström et al. (2024) phylogenetic trees. For each language, we retrieve the path from its family root node to the language node. The relatedness of two languages (gen) is the ratio of shared nodes along these paths. [^3] Because different trees/branches have different depths (owing to different family sizes and/or different levels of documentation), comparisons of scores can only serve as a proxy for the degree of phylogenetic relatedness. This is also noted in the metadata for lang2vec, which calculates genetic distances in a similar way: http://www.cs.cmu.edu/~dmortens/projects/7_project/. For constructed languages, we ignore the top-most level (e.g., “artificial languages”) and instead retrieve phylogenetic information starting at the second level.
@@
We use lang2vec’s location information (each language is represented as a vector of distances to a number of points on the Earth’s surface) and calculate the Euclidean distances between language vectors. [^4] We do not use lang2vec’s precomputed distances, as the package erroneously returns the maximum distance for some languages, regardless of the comparison. We define geographic proximity (geo) as 1 minus the distance.
@@
Correlations between similarity measures can influence how importance is assigned to different measures in regression analyses. Table 8 in §B shows how the different similarity measures are correlated with each other. Lexical similarity and phylogenetic relatedness are highly correlated ($r=0.87$). Additionally, the two grammar-based similarity measures (gb and syn) are moderately strongly correlated ($r=0.61$), as are gb and lexical/phylogenetic similarity ($r=0.57$, $0.56$), and word overlap in UD and lexical similarity ($r=0.59$).
@@ 4 Supporting Results
Here we present the results on which the main analyses in §5 build. In §4.1, we present the POS tagging, parsing and topic classification scores within and across languages. In §4.2, we compare how (dis)similar transfer patterns are across tasks and experiments, which helps contextualize the importance of different similarity measures for different experiments and tasks in §5.1.
@@
As expected, the within-language performances are much higher than the cross-lingual performances. Table 2 shows the mean scores within and across languages and datasets. In the following analyses, we focus on the language-level results. For UD, which has nearly twice as many datasets as languages, we find that trends are very stable across datasets of the same language (§D.3).
All of our models achieve reasonable within-language results, indicating that cross-lingual transfer (to an appropriate target language) could be feasible for all models. For the two straightforward classification tasks (POS tagging and topic classification), we construct random baselines (§D.1) that are outperformed by all models in within-language evaluations. For our topic classification models, the within-dataset performance is comparable to the baselines in the dataset paper (Adelani et al., 2024), but we do not match the performance of their best model. Our POS tagging results are similar but not identical to the ones that de Vries et al. (2022) obtained on UD v2.8 with fine-tuned XLM-R models for 65 training and 105 test languages. If we consider our own POS tagging results but only select the subset of languages that was used in de de Vries et al.’s experiments, the POS tagging accuracies are highly correlated with those that de Vries et al. obtained (Pearson’s $r$ and Spearman’s $\rho=0.73$, $p<0.0001$). Thus while the model choice (UDPipe vs. XLM-R) matters for the resulting patterns, many transfer trends are similar.
Dataset level Language level Within Across Within  Across Grammatical tasks POS accuracy 96.4 3.1 43.6 20.8 93.4 6.8 43.0 20.2 POS acc. (de Vries et al.) 94.1* 4.5 57.4* 22.4 94.1 4.5 57.4 22.4 UAS 88.7 6.4 37.3 19.3 84.8 10.1 36.8 18.8 LAS 84.6 8.6 21.2 17.9 78.4 13.7 20.6 17.0 Topic classification (accuracy) topics-base 70.3 4.0 20.7 8.8 66.7 14.1 20.1 8.8 topics-translit 69.4 4.1 22.5 8.2 66.7 10.8 22.5 8.2 topics-mbert 60.1 18.7 42.6 20.4 58.2 19.9 42.6 20.4 MLP (Adelani et al.) 62.3 XLM-RB (Adelani et al.) 70.9 XLM-RL (Adelani et al.) 76.1
Table 2: Scores of models evaluation within vs. across datasets and languages, in this work and related work. Our analyses focus on the language-level scores. Scores are mean scores (in %), with standard deviations in subscripts. XLM-RB and L = XLM-R base/large. *When multiple datasets were available for one language, de Vries et al. combined them into one.
@@
Although topics-base and topics-translit achieve the same within-language accuracy, topics-translit performs slightly better cross-lingually, having the advantage of higher $n$-gram overlap between datasets. [^6] We further analyze the effect of writing systems in §D.4. In our correlation analysis (§5.1), (lack of) character overlap serves as a proxy for writing system differences. In cross-lingual evaluations, topics-mbert benefits from its multilingual input embeddings: its accuracy is about twice as high (42.6%) as for the monolingual models. Generally, transfer works well if both training and test languages are in mBERT’s pre-training data (or closely related to a language that is): The average accuracy is 68.3% if both languages were included in mBERT’s pre-training data, 46.5% if only the target language was included, 36.1% if only the source language was included, and 33.2% if neither was included. This is consistent with results by Adelani et al. (2024), who fine-tuned four XLM-Rlarge models on one high-resource language each and found their transfer results to be very similar to each other.
@@
Figure 2: Different experiments produce different transfer patterns. NLP transfer results for all combinations of training (columns) and test languages (rows). The darker a cell, the higher the score. The three heatmaps for the grammatical tasks are sorted in the same order, and the three heatmaps with the topic classification results are sorted in the same order. The darker diagonal shows the within-language scores. Large, labelled heatmaps are in Appendix §D.2.
@@ 4.2 Comparing transfer patterns across tasks
In Figure 2, we plot heatmaps of the transfer results for the different experiments. Darker colours correspond to higher scores. To save space, we omit the language labels (which can be found in §D.2). We compute correlations between the results, which we summarize below (details in §E.1).
@@ 5.1 How do the similarity measures correlate with the tasks and input representations?
We calculate the correlations between the similarity measures and the NLP task results. To account for differences between test sets, we calculate correlations (Pearson’s r) between similarity measures and transfer results on a test language level. We compare the performance of different NLP systems on the same test language, but we do not compare the performance of a single system across multiple test sets. This decision is motivated especially by the challenges of comparing parsing performance across treebanks: Sentence length influences parsing difficulty as it determines the available search space (cf. McDonald and Nivre, 2011, Choi et al., 2015), and morphological differences between languages hamper the comparability of parsers Nivre and Fang (2017).
We analyze overall correlation scores by averaging across test languages. We treat correlation coefficients with p-values of at least 0.05 as 0, and exclude items where we could not calculate similarity scores due to missing linguistic information. We use language-level averages so that those languages with multiple training and test datasets do not have artificially high correlations, and languages with multiple test datasets do not have greater influence on the overall averages. However, if we instead consider treebank-level correlations, we observe that test sets that belong to the same language show very similar correlation patterns and scores (§D.3).
The correlations between task results and similarity measures vary across our experiments. Figure 3 shows the correlations, which we summarize below. Unaggregated correlations and 95% confidence intervals of the mean scores can be found in §E.3. These unaggregated correlations show that, while there are some outliers, the overall trends we report below accurately reflect the language-level trends.
@@
We observe some common trends across experiments: Training dataset size does not matter much: it is the same across all topic classification experiments, and the correlations between training set size and performance in the grammatical tasks are close to zero. The phonological and phonetic similarity measures (pho, inv) also generally have low correlation scores. [^8] The correlations with inv are comparatively high for the $n$-gram-based topic classification models. We assume that this is due to inv being moderately correlated with string similarity measures like trigram overlap (§B), which is in turn strongly correlated with the topic classification results.
The strongest predictor for parsing performance is syntactic similarity (syn) as determined by lang2vec ($r$avg_syn$=0.57$), followed by the similarity of Grambank features (gb; $r$avg_gb$=0.42$). These correlations are even stronger when we only consider the test languages that mBERT was pretrained on ($r$avg_syn$=0.69$ and $r$avg_gb$=0.60$, respectively; Appendix §E.3).
@@
The transfer results cannot be predicted by any one factor alone. We fit a linear mixed effects model for each experiment, with the NLP score as the dependent variable and the source and target languages as random effects. The fixed effects are the similarity scores, whether the training and test data use the same writing systems, and whether the test language was included in mBERT’s pre-training data. We use R R Core Team (2024) and the lme4 package Bates et al. (2015). This analysis involves a reduced set of languages, as for many language pairs in our data at least one similarity metric is not defined and we do not perform any data imputation. The analysis shows similar trends to the correlations described above. Additionally, for each experiment, most of the variables in our analyses are significant predictors of the transfer score (details in §E.2). This even sometimes applies to measures that capture similarity at the same linguistic level: e.g., both grammatical similarity measures (syn, gb) are independently among the highest predictors for parsing performance. The finding that multiple measures contribute to predicting the transfer score confirms previous work Lin et al. (2019); de Vries et al. (2022).
@@ 5.2 Practical implications
size pho inv geo syn gb gen lex chr word* Top-1 source candidate (= most similar language) POS deV 88 1011 1011 78 57 811 68 POS 2912 1513 1412 1515 1012 1212 0911 1010 1514 1213 LAS 2114 1312 1313 1316 710 1009 810 809 1616 1113 UAS 2712 1613 1512 1615 809 1311 910 1011 1714 1415 top.-b. 1712 1714 1314 1512 1413 912 911 1311 405 top.-tr. 1310 1311 $1$ 1109 1009 709 708 2013 304 top.-m. 1213 1113 1011 910 809 809 809 1210 913 Top-3 source candidates POS deV 34 45 45 25 24 58 24 POS 2513 708 708 507 304 406 507 406 708 507 LAS 1814 809 709 508 304 406 406 305 811 609 UAS 2413 808 708 608 305 506 507 507 910 709 top.-b. 1010 911 509 607 609 507 306 809 204 top.-tr. 808 709 408 506 506 305 305 1411 103 top.-m. 607 505 405 404 405 404 405 504 611
Table 3: Mean performance loss if picking source languages based solely on a given metric (or based on training dataset size). Performance loss in percentage points. Subscripts = standard deviations; “word*” = wor, tri, swt; deV = de Vries et al. (2022).